OpenStudy (rational):

Evaluate \(\sum_{n=0}^\infty x^{n} (n^{2} + n)\) assuming the series converges

OpenStudy (amistre64):

x has to be between 0 and 1 then ..... right?

OpenStudy (rational):

yes.. |x| < 1 for convergence

OpenStudy (amistre64):

and a power function (not technical term for it) grows faster than a poly right? so it would at least converge

OpenStudy (amistre64):

0r^0 + 0r^0 1r^1 + 1r^1 4r^2 + 2r^2 9r^2 + 3r^3 the right side we did earlier converges to 1/(1-r)^2

OpenStudy (rational):

Ahh we can use that i guess earlier we did \(\sum\limits_{n=0}^{\infty} nx^n = \dfrac{x}{(1-x)^2}\)

OpenStudy (rational):

*x in numerator

OpenStudy (amistre64):

. = 1 + r + 4r^2 +9r^3 +16r^4 + ... + (n-1)^2 r^(n-1) -r. = - r - r^2 - 4r^3 - 9r^4 - ... - (n-2)^2 r^(n-1) -(n-1)^2r^n ----------------------------------------------------------- (1-r).= 1 +3r^2 +5r^3 +7r^4 + .... + (2n-3) r^(n-1) - (n-1)^2 r^n and the next one of course reduces to constants

OpenStudy (amistre64):

(1-r).= 1 +3r^2 +5r^3 +7r^4 + .... + (2n-3) r^(n-1) - (n-1)^2 r^n -r(1-r).= -r -3r^3 -5r^4 - .... - (2n-3) r^n + (n-1)^2 r^(n+1) --------------------------------------------------------------- (1-r)^2, = 1-r+3r^2 +2r^3 +2r^4 + ... + 2r^(n-1) -[(n^2-2] r^n + (n-1)^2 r^(n+1) anohter -r seems to be able to get us there. with just a few term left over

OpenStudy (rational):

that works nicely, the sum n^2r^n doesn't start with 1 ``` . = r + 4r^2 +9r^3 +16r^4 + ... + (n-1)^2 r^(n-1) -r. = - r - r^2 - 4r^3 - 9r^4 - ... - (n-2)^2 r^(n-1) -(n-1)^2r^n ----------------------------------------------------------- (1-r).= 3r^2 +5r^3 +7r^4 + .... + (2n-3) r^(n-1) - (n-1)^2 r^n ``` and the next one of course reduces to constants

OpenStudy (amistre64):

n^2 r^n by the wolf gets us reduced to r^2+r/(1-r)^3 but id have to work it on paper to keep it organized better :)

OpenStudy (rational):

nooo, what we had will do haha we're mostly there :)

OpenStudy (amistre64):

0r^0=0 , yeah :)

OpenStudy (rational):

x = r in this thread

OpenStudy (rational):

*then SUM x^n (n^2+n) evaluates to (r^2+r)/(1-r)^3 + r/(1-r)^2 = 2r/(1-r)^3

OpenStudy (amistre64):

thought right term was 1/(1-r)^2 or am i picking a typo?

OpenStudy (rational):

\[\dfrac{1}{1-r} = \sum\limits_{n=0}^{\infty} r^n \] differentiating both sides gives \[\dfrac{1}{(1-r)^2} = \sum\limits_{n=0}^{\infty} nr^{\color{red}{n-1}} \] we need to multiply \(r\) both sides to fix that exponent on right hand side right

OpenStudy (amistre64):

right, got it :)

OpenStudy (amistre64):

ive always wondered if it was just convention, but derivative of a summation index adds 1

OpenStudy (amistre64):

\[y=\sum_0a_nx^n\] \[y'=\sum_1a_nnx^{n-1}\]

OpenStudy (rational):

Ahh right but there is no harm in starting index at 0 right, we will just get the first term 0 or will it shift the \(a_i\) 's

OpenStudy (rational):

\[\sum\limits_{n=0}^{\infty} nr^{n} = \sum\limits_{n=1}^{\infty} nr^{n} \]

OpenStudy (amistre64):

that seems fair for the first derivative :)

OpenStudy (rational):

we get more leading 0's as we differentiate more times other than that.. hmm is there any harm

OpenStudy (amistre64):

kx^0 if we just powe rule it is: 0k/x = 0

OpenStudy (rational):

i see.. that 0 coefficient still makes it 0

OpenStudy (rational):

but using power rule on x^0 doesn't look correct in series yeah

OpenStudy (rational):

that might be the reason we explicitly change the start of index when differentiating hmm

OpenStudy (amistre64):

something like that yes. helps avoid some technical issues.

OpenStudy (amistre64):

\[\frac{a_0}{1-r}=\sum_0^{inf}a_n r^n\] \[\frac{a_0}{1-r}=a_0+\sum_1^{inf}a_n r^n\] \[-\frac{a_0}{(1-r)^2}=0+\sum_1^{inf}a_nn r^{n-1}\]

OpenStudy (amistre64):

\[-\frac{a_0}{(1-r)^2}=\sum_0^{inf}a_{n+1}(n+1) r^{n}\]

OpenStudy (amistre64):

do i have an error?

OpenStudy (amistre64):

the negative is an error ... that gets me everytime ..

OpenStudy (rational):

i think \(a_n \) needs to be a constant for using that infinite geometric series formula a1/(1-r)

OpenStudy (rational):

\[\frac{a_0}{1-r}=\sum_0^{inf}a_n r^n\] this works oly if \(a_n = a_0\) for all \(n\)

OpenStudy (amistre64):

true true, but at 315am it all looks the same :)

OpenStudy (rational):

lmao i have another way to work this problem but may be il post it later as i don't want to keep u from sleeping lol

OpenStudy (amistre64):

very well you have a good night/morning wherever you are ;)

OpenStudy (rational):

gnite!